# DCoM: Active Learning for All Learners

### v2

@@ C.3 Active sampling
Time complexity: Active sampling involves balancing model min-margin and example density based on the current adjacency graph. Initially, the margin of each example is computed using the softmax output from the model’s last layer, requiring $O(n)$ time. Subsequently, the current adjacency is created using $\delta_{\text{avg}}$, which is the average over the ${\mathbb{L}}$ $\delta$ values list - $\Delta$. As discussed in App. C.1, this step has a time complexity of $O(dn^{2})$. Following this, the iterative process for selecting a single sample includes the following steps:
@@
Appendix D Additional empirical results
D.1 $\delta_{0}\,\,\,$ Initialization
To set the initial value $\delta_{0}$, we adopted the method outlined in [39], as detailed in Section 3. Fig. 7 displays the purity function across various $\delta$ values, along with the selected $\delta$ values for each dataset. As mentioned previously, the $\delta$ values for CIFAR-10 and CIFAR-100 are derived from [39].
@@ D.1 $\delta_{0}\,\,\,$ Initialization
D.2 $\Delta\,\,\,\,$ distribution throughout the training process
@@
(c) ImageNet-50
Figure 8: The $\Delta$ distribution through DCoM algorithm. Each plot displays the mean and standard error over 5-10 repetitions. The results indicate that as the emphasis shifts towards maximizing the purity of sample balls, rather than solely focusing on maximum coverage, the radii adapt accordingly. This, combined with the improved performance observed with DCoM, highlights the necessity of the dynamic algorithm in enhancing the effectiveness of active sampling.
D.3 Dynamic algorithm coverage
The analysis presented in Fig. 9 demonstrates a consistent behavior of probability coverage across different datasets during the DCoM process. This indicates that using probability coverage as a measure of progress consistently proves to be a wise decision.
(a) CIFAR-10 and STL-10
(b) CIFAR-100 and ImageNet-50
Figure 9: Comparison between $P(C({\mathbb{L}},\Delta))$ values during DCoM sampling process, across several datasets. Each plot displays the mean and standard error over $3-10$ repetitions. These figures demonstrates a consistent behavior of probability coverage.
D.4 Decomposing the algorithm
In Section 4.3, we present an ablation study where we evaluate the contribution of each component in objective function (1) separately on CIFAR-100 dataset. In Fig. 10 you can see the results over CIFAR-100 and ImageNet-50 datasets.
(a) CIFAR-100
(b) ImageNet-50
Figure 10: Performance evaluation (mean accuracy and STE) when optimizing each component of objective function (1) separately and together. DPC denotes the Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM. Margin is DCoM’s choice for ${{\mathcal{O}_{high}}}(x)$.
In Fig. 11 you can see the same ablation but using entropy as ${{\mathcal{O}_{high}}}(x)$ instead of margin.
(a) CIFAR-10
(b) CIFAR-100
(c) ImageNet-50
Figure 11: Assessment of each algorithm’s objective function individually across multiple datasets. DPC stands for Dynamic Probability Coverage, corresponding to ${{\mathcal{O}_{low}}}(x)$ in DCoM. DCEM represents DCoM with entropy as ${{\mathcal{O}_{high}}}(x)$ instead of margin. Each plot displays the mean and standard error over $5-10$ repetitions. These graphs demonstrate how the weighted objective function consistently produces superior results across all labeled set sizes $|{\mathbb{L}}|$.
D.5 Different embedding space
In Section 4.3, we present an ablation study where we repeat the basic fully-supervised experiments while varying the embedding employed by DCoM. In Fig. 12 you can see the same ablation using BYOL [13] representation.
D.6 Comparison per dataset
Tables 5-6 present the empirical accuracy values of several datasets using different active learning algorithms. In each active step, a new learner is trained from scratch over all available labeled data, and the results are presented as the [mean $\pm$ STE] over $5-10$ repetitions (3 for ImageNet subsets). As mentioned earlier, the results for LDM-s algorithm were acquired from the study by [4]. Their lack of code provision suggests discrepancies in the running setup. Nonetheless, the disparity between their method and random selection in their setup is smaller than that observed in ours, implying the superiority of our method. The table includes the results for all datasets, with the name of each dataset below it.
Table 5: Model accuracy for different sizes of labeled set ${\mathbb{L}}$ and AL strategies
@@ D.6 Comparison per dataset
Table 6: Model accuracy for different sizes of labeled set ${\mathbb{L}}$ and AL strategies

### v3

@@ D.3 Active sampling
Time complexity: Active sampling involves balancing model min-margin and example density based on the current adjacency graph. Initially, the margin of each example is computed using the softmax output from the model’s last layer, requiring $O(n)$ time. Subsequently, the current adjacency is created using $\delta_{\text{avg}}$, which is the average over the $\mathbb{L}$} $\delta$ values list - $\Delta$. As discussed in Appendix D.1, this step has a time complexity of $O(dn^{2})$. Following this, the iterative process for selecting a single sample includes the following steps:
@@
Appendix E Additional empirical results
E.1 $\boldsymbol{\delta_{0}}$ Initialization
To set the initial value $\delta_{0}$, we adopted the method outlined in (Yehuda et al. 2022), as detailed in Section 3. Fig. 9 displays the purity function across various $\delta$ values, along with the selected $\delta$ values for each dataset. As mentioned previously, the $\delta$ values for CIFAR-10 and CIFAR-100 are derived from (Yehuda et al. 2022).
@@ E.1 $\boldsymbol{\delta_{0}}$ Initialization
E.2 $\boldsymbol{\Delta}$ distribution evolution throughout training
@@
Figure 10: The $\Delta$ distribution through DCoM algorithm. Each plot displays the distribution over 5-10 repetitions. The results indicate that as the emphasis shifts towards maximizing the purity of sample balls, rather than solely focusing on maximum coverage, the radii adapt accordingly. This, combined with the improved performance observed with DCoM, highlights the necessity of the dynamic algorithm in enhancing the effectiveness of active sampling.
@@
Table 3: Model accuracy for different sizes of labeled set $\mathbb{L}$} and AL strategies.
@@
Table 4: Model accuracy for different sizes of labeled set $\mathbb{L}$} and AL strategies.
[…]
Fig. 11 tracks the behavior of the competence score $S_{\mathcal{L}}$} across 3 datasets. In all cases, the comparison of Figs.11(a),11(b) reveals that $S_{\mathcal{L}}$} increases monotonically with competence as measured by accuracy. Interestingly, while the critical point where $S_{\mathcal{L}}$} changes from favoring $\mathcal{O}_{low}$} (small values) to favoring $\mathcal{O}_{high}$} (large values) corresponds to an intermediate level of accuracy, this transition is delayed for the more difficult datasets.
(a) Accuracy with random sampling.
(b) $S_{\mathcal{L}}$} during DCoM training.
Figure 11: (a) Accuracy (mean and standard error) as a function of budget $b$, learning to classify CIFAR-100, ImageNet-50, and ImageNet-100 using a random sample of $b$ points. (b) The competence score $S_{\mathcal{L}}$} as a function of budget.
E.4 Different embedding space
In Section 4.3, we present an ablation study where we repeat the basic fully-supervised experiments while varying the embedding employed by DCoM. In Fig. 12 you can see the same ablation using BYOL (Grill et al. 2020) representation.
E.5 Comparison per dataset
Tables 3-4 present the empirical accuracy values of several datasets using different active learning algorithms. In each active step, a new learner is trained from scratch over all available labeled data, and the results are presented as the [mean $\pm$ STE] over 5105-10 repetitions (3 for ImageNet subsets). As mentioned earlier, the results for LDM-s algorithm were acquired from the study by (Cho et al. 2023). Their lack of code provision suggests discrepancies in the running setup. Nonetheless, the disparity between their method and random selection in their setup is smaller than that observed in ours, implying the superiority of our method. The table includes the results for all datasets, with the name of each dataset below it.